# Allocating Positional Goods: A Mechanism Design Approachthanks: I thank Krishna Dasaratha, Alex Gershkov, Bart Lipman, Chiara Margaria, and Juan Ortner for their helpful comments. The paper subsumes results in my note “A Mechanism Design Approach to ‘Gainers and Losers in Priority Services’.”

### v2

@@ Abstract
I study the optimal allocation of positional goods in the presence of externalities arising from consumers’ concerns for relative consumption. Using a mechanism design approach, I characterize the externalities by a feasibility condition. I show that the revenue-maximizing mechanism possibly excludes some buyers and fully separates the participants if their type distribution satisfies Myerson’s regularity. The seller can guarantee at least half the maximum revenue by offering a single level of positional goods. When there is no exclusion, as the seller offers more positional good levels, the consumer surplus decreases (increases) if the distribution has an increasing (decreasing) failure rate. Higher participation increases the consumer surplus under increasing failure rates. Applications include luxury goods, priority services, education, and organizational hierarchies.
Keywords: Positional goods, mechanism design, externalities, feasibility.
@@ 1 Introduction
Positional goods are goods or services whose value to consumers depends on their relative positions in consumption. Hirsch (1976) first coined the concept of positional good that is “either scarce in some absolute or socially imposed sense” or “subject to congestion or crowding through more extensive use” (pp. 27). Because consumers care about their relative positions, allocating an item to one buyer has externalities on others.
@@
Examples of positional goods abound. Consumer goods, especially luxury goods, are often positional because people derive happiness from comparisons (Frank, 1985a; Luttmer, 2005; Carlsson et al., 2007). Another example is priority services and front-running (Gershkov and Winter, 2023), such as priority boarding. Such services reduce consumers’ waiting time by moving them ahead of the crowd, but they become less effective as more consumers use them. Additionally, education is the classic example in Hirsch (1976): “The value to me of my education depends not only on how much I have but also on how much the man ahead of me in the job line has.”
More broadly, status within an organization can also be viewed as a positional good, as employees invest efforts to move up in the hierarchy (Moldovanu, Sela, and Shi, 2007). As a revenue-maximizing seller, the organization designer can also induce effort through employees’ positional concerns and exclude low-ability employees.
@@
In these examples, the allocation of positional goods has externalities on consumers, as moving up the position of one consumer would push down that of another. At the same time, when more consumers buy the same good (e.g., priority boarding), its value depreciates due to more extensive use. Thus, unlike ordinary goods, it is impossible to allocate the good (e.g., the first place in line) to multiple consumers at face value: “What each of us can achieve, all cannot” (Hirsch, 1976). In general, positional goods induce a game of status, in which agents pay in money or effort to attain higher relative positions. 111 Because of the dual interpretation of the model, I use “consumers” and “agents”, as well as “seller” and “principal” interchangeably. While the principal can exploit the status game to incentivize spending or effort, this often comes at the expense of agent welfare. 222 For example, Asian governments have been orchestrating educational disarmament (The Economist, 2021). U.S. colleges switched to coarse grading, such as pass/fail, to relieve student stress during the COVID-19 pandemic. Luxury consumption is discouraged by many cultures and religions.
@@
In this paper, I study the optimal allocation of positional goods using a mechanism design approach. I assume buyers care about their relative positions, or status, which depends on the mass of consumers who buy goods at the same level and at lower levels. Moreover, they have private information about their types—i.e., valuations of the status (or marginal costs of effort). The seller can offer multiple levels of positional goods at different prices. The buyer can pay a price (or effort) for one of the goods to obtain the status value, or pay nothing and opt out.
@@
First, I characterize positional externalities by a feasibility condition: the interim status allocation is feasible (i.e., can be induced by an allocation of positional good) if and only if it is weakly majorized by the buyer’s type distribution in the quantile space.
@@
Then, I focus on revenue maximization. Using results in Kleiner, Moldovanu, and Strack (2021, henceforth KMS), I show that offering more levels of positional goods increases the seller’s revenue if the type distribution is regular in the sense of Myerson (1981). Therefore, the revenue-maximizing mechanism excludes types with negative marginal revenue and fully separates participants, which can be implemented by an all-pay auction with a reserve price, where the buyer who pays more obtains the higher position. In contrast, if Myerson’s regularity condition fails for some types, the optimal mechanism pools them into the same level of positional goods.
I also find that a simple posted-price mechanism that offers a single level of positional goods can guarantee at least half the maximal revenue. In the proof, I consider an auxiliary screening problem of selling a continuum of ordinary (i.e., nonpositional) goods to a continuum of buyers. The screening problem highlights the distinction between selling positional and ordinary goods: in the former, allocating the good to one buyer has externalities on others, so the seller cannot sell the highest status to multiple buyers at face value. Therefore, selling positional goods results in lower revenue compared to selling ordinary goods (Mussa and Rosen, 1978).
@@
Next, I study how the number of positional good levels and exclusion affect consumer welfare. I first focus on the case when consumer participation is fixed, for example, when there is a free regular service available. I show that the consumer surplus is higher when the seller offers more positional good levels if the buyer’s type distribution has an increasing failure rate (IFR), which implies a thin tail. This result highlights the tension between the seller’s revenue and the consumer surplus. On the other hand, when the type distribution has a decreasing failure rate (DFR), which implies a heavy tail, the consumer surplus is higher when the seller offers fewer positional good levels. Furthermore, under IFR, expanding consumer participation increases the consumer surplus, as the marginal gain of additional participants outweighs the inframarginal loss of the existing participants from reduced status due to pooling.
@@
I also study optimal mechanisms that maximize the consumer surplus. When negative transfers are allowed, full separation without exclusion maximizes the consumer surplus. Intuitively, this is because assortative matching, and possibly negative transfers can redistribute the efficiency gains among consumers (see also Gershkov and Schweinzer, 2010). By contrast, in applications where the price (effort) must be nonnegative, such as organizational hierarchies and education, I show that total pooling without exclusion maximizes the consumer surplus under IFR. Unlike selling ordinary goods, exclusion raises the status of participants who would be otherwise pooled with the excluded ones. Nevertheless, under IFR, the harm of exclusion to the excluded consumers exceeds its benefit to the participants. In contrast, if the distribution has a decreasing failure rate or a heavy tail (e.g., the Pareto distribution), the consumer surplus under the nonnegative price constraint is maximized by full separation without exclusion.
@@
As an extension, I consider the case where the intrinsic value of the positional good is determined by its intrinsic quality produced at a cost (à la Mussa and Rosen (1978)), and the results remain robust. Moreover, I also allow the seller to deliberately lower the status, or equivalently, prolong the waiting time by pausing service or procrastinating. Thus, the seller can create negative status or excessive waiting time as a substitute for exclusion. I show that the revenue-maximizing mechanism fully separates those with positive virtual valuations and assigns negative status to those who with negative virtual valuations. Additionally, I consider cases where the status game is a “suffering” in that higher types suffer more from participation, and the results remain robust.
@@
The model has dual interpretations that lead to two strands of applications. On the one hand, it is a monopolist problem of selling positional goods, such as priority services (Gershkov and Winter, 2023) and luxury goods. Alternatively, it can be interpreted as an organizational design problem where the principal motivates agent effort (e.g., Moldovanu, Sela, and Shi, 2007; Immorlica, Stoddard, and Syrgkanis, 2015) or maximizes agent welfare through status within an organization. Under this interpretation, the optimal mechanism that maximizes effort possibly excludes some types and fully separates the participants if the ability distribution satisfies Myerson’s regularity. This can be seen in “rank-and-yank” organizations that rank every employee and terminate those who underperform.
@@
The implication also echoes the education example by Hirsch (1976) and sheds light on the consequences of education arms races. From the agent’s perspective, while the common wisdom calls for educational disarmament and less intense competition, I show that this benefits agents if their ability distribution has a thin tail. Under the IFR property, admission lotteries or random assignments, instead of meritocratic allocation, maximizes agent welfare. On the contrary, if the ability distribution has a heavy tail, meritocracy maximizes agents’ welfare (and their effort). Additionally, if education is sufficiently socially productive, educational arms races can be optimal for social welfare.
@@
In Appendix A, I consider the case where exclusion is impossible, which is true in many applications where the lowest status is free. For example, a society may not allow exclusion, so investing no effort guarantees the lowest social status (Moldovanu, Sela, and Shi, 2007). Moreover, priority services may be offered in addition to free regular services (Gershkov and Winter, 2023). In these cases, the lowest-level positional good may still have a positive value, as the lowest status cannot be offered to multiple agents at face value either. I provide the necessary and sufficient conditions for the revenue-maximizing mechanisms to be separating or pooling and for consumer welfare to be decreasing or increasing in the number of positional good levels, as well as some additional results, such as approximation by offering two levels. The results are qualitatively similar to the benchmark model with exclusion.
@@
The idea that the consumer’s utility depends on the comparison with others’ consumption dates back to Veblen (1899) (see also Duesenberry, 1949; Leibenstein, 1950; Bagwell and Bernheim, 1996). Since Hirsch (1976) introduced the concept of positional good, the status interdependence in preference has been formally modeled and studied by Frank (1985b, 2005, 2008), Robson (1992), and Hopkins and Kornienko (2004).333 Another strand of literature uses the Duesenberry formulation that stems from Duesenberry (1949) and Pollak (1976) (see the survey by Truyts (2010)), which assumes the consumer’s utility depends on the consumption of others in addition to their absolute consumption. This literature focuses on consumer choice in the presence of positional externalities and the welfare effects of changes in the income distribution and government policies (e.g., luxury tax).
@@
Empirical evidence of positional goods and externalities include Frank (1985a, 2005), Luttmer (2005), Alpizar et al. (2005), Carlsson et al. (2007), Bursztyn et al. (2018). In particular, education is a typical example of positional good (Hirsch, 1976). Durst (2021) finds empirical evidence that education is indeed positional in Germany.
@@
Moldovanu, Sela, and Shi (2007, henceforth MSS) study the optimal design of the number and size of status categories in an organization to maximize agents’ effort when they care about their relative position (i.e., status). They find that full separation is optimal if the ability distribution satisfies IFR, while coarse partition may be optimal depending on the distribution—for example, the optimal partition involves only two categories if it is sufficiently concave. Despite the similarities in our results, this paper is different in three ways.444 The definition of status in MSS, which equals the number of strictly inferior agents minus the number of strictly superior agents, is equivalent (up to an affine transformation) to my specification (see Footnote 10). First, I allow for exclusion of agents and stochastic allocations and study agent welfare. In reality, the exclusion of agents is common in organizations, and employee welfare is an important consideration. Second, in the case without exclusion, I obtain stronger results that are either necessary and sufficient conditions or require fewer distributional assumptions (see Appendix A). Third, I use a mechanism design approach with a feasibility characterization using majorization, while they use methods in optimal contest design (with finite agents).
@@
Relatedly, Rayo (2013) studies a model of monopolist provision of conspicuous goods and characterizes the optimal allocation. His results are qualitatively similar to MSS’s, and he does not consider exclusion, consumer welfare maximization, or stochastic mechanisms either. Conspicuous goods in his definition also differ from positional goods (or status in MSS) in that consumers care about the average type in their assigned category instead of relative positions. Nevertheless, the feasibility condition can also be applied to conspicuous goods in his model (see Remark 4), which explains the similarities between his results and those of positional goods.555 Board (2009) assumes more general benefits from being in a category (i.e., group quality) and finds a contradictory result that the revenue-maximizing principal segregates agents too finely compared to the efficient level (which is full separation in my model). He also finds that revenue maximization leads to too much exclusion.
@@
Similar to MSS, Immorlica, Stoddard, and Syrgkanis (2015) study the optimal design of badges to incentivize users’ contribution (i.e., effort) to the online community through the status value. Unlike MSS, they consider exclusion and allow the status value to be concave or convex in the relative position. They find the “leaderboard with a cutoff” mechanism is optimal if the distribution satisfies Myerson’s regularity. As in MSS, they do not consider agent welfare or stochastic allocations, which is an important consideration in online communities. Most importantly, their status value function is strictly decreasing in the number of people weakly superior to them (without adjustment to those at the same level), which naturally precludes pooling from being optimal because it decreases the status of every agent pooled together. Consequently, (i) total pooling will never be optimal—even for agent welfare—in sharp contrast to my result, (ii) the optimality of full separation in effort maximization is less surprising, and (iii) exclusion is equivalent to pooling at the lowest status.
@@
In a different vein, Gershkov and Winter (2023, henceforth GW) study a model of priority service and analyze its implications for consumer welfare. Priority service (PS) is an example of positional goods that arise from technological constraints. Using a mechanism design approach, this paper generalizes their model by allowing for multiple (and stochastic) PS levels and exclusion. By contrast, they focus on two PS levels instead of the optimal mechanism. 666 They extend their model to multiple PS levels but without exclusion in Section V. In the case without exclusion, I refine some of their results by providing necessary and sufficient conditions and show some additional results in Appendix A.
@@
Another closely related paper is by Loertscher and Muir (2022), who study revenue maximization when selling a fixed inventory of goods with heterogeneous qualities. Without the intrinsic value and under linear status, my model is equivalent to a continuous version of theirs where each quality has a unit stock. In particular, pooling multiple types into the same status corresponds to randomization (or conflation) in their model.
@@
More broadly, this paper is also related to mechanism design with allocative externalities (e.g. Jehiel, Moldovanu, and Stacchetti, 1996; Jehiel and Moldovanu, 2005; Dworczak, Reuter, Kominers, and Lee, 2024).

### v3

@@ Abstract
I study the optimal allocation of positional goods in the presence of externalities arising from consumers’ concerns about relative consumption. Applications include luxury goods, priority services, education, and organizational hierarchies. Using a mechanism design approach, I characterize the set of feasible allocations through a majorization condition. The revenue-maximizing mechanism possibly excludes some buyers and fully separates participants under Myerson’s regularity condition. The seller can guarantee at least half the maximum revenue by offering a single good. Without exclusion, offering more levels of goods decreases (increases) consumer surplus under increasing (decreasing) failure rates. Higher participation raises consumer surplus under increasing failure rates.
Keywords: Positional goods, mechanism design, externalities.
@@ 1 Introduction
Positional goods are goods or services whose value to consumers depends on their relative positions in consumption. Hirsch (1976) first coined the concept of positional good that is “either scarce in some absolute or socially imposed sense” or “subject to congestion or crowding through more extensive use” (pp. 27). In essence, because consumers care about their relative positions, the allocation of positional goods creates externalities on others.
@@
Examples of positional goods abound. Consumer goods are often positional because people derive happiness from comparisons (Frank, 1985a; Luttmer, 2005; Carlsson, Johansson-Stenman and Martinsson, 2007). Another example is priority services and front-running (Gershkov and Winter, 2023), such as priority boarding. These services reduce consumers’ waiting time by moving them ahead of the crowd, but they become less effective as more consumers use them. Additionally, education is a classic example in Hirsch (1976): “The value to me of my education depends not only on how much I have but also on how much the man ahead of me in the job line has.” More broadly, status within an organization can be viewed as a positional good, as employees invest effort to move up in the hierarchy (Moldovanu, Sela and Shi, 2007). Similar to a revenue-maximizing seller, the organization designer induces effort through employees’ positional concerns and by excluding low-ability employees.
@@
In these examples, the allocation of positional goods has externalities on consumers. On the one hand, moving up the position of one consumer would push down that of another. On the other, when more consumers buy the same good, its value depreciates due to more extensive use. Therefore, unlike ordinary goods, it is impossible to allocate the same positional good (e.g., the first place in line) to multiple consumers without reducing its value: “What each of us can achieve, all cannot” (Hirsch, 1976). For this reason, positional goods remain scarce and competitive despite economic growth. In general, positional goods induce a game of status, in which agents pay money or exert effort to attain higher relative positions.  [^1] Because the model has implications for both problems, I use “consumers” and “agents” interchangeably, as well as “seller” and “principal.” While the principal can exploit the status game to incentivize spending or effort, this often comes at the expense of agents’ welfare.  [^2] For example, Asian governments have been orchestrating educational disarmament (The Economist, 2021). U.S. colleges switched to coarse grading, such as pass/fail, to relieve student stress during the COVID-19 pandemic. Luxury consumption is discouraged by many cultures and religions.
@@
In this paper, I study the optimal allocation of positional goods using a mechanism design approach. I assume buyers care about their relative positions, or status, which depends on the mass of consumers who buy goods at the same or lower levels. Moreover, they have private information about their types—i.e., valuations of the status (or marginal costs of effort). The seller can offer multiple levels of positional goods at different prices. The buyer can pay a price (or exert effort) for one of the goods to attain the status, or pay nothing and opt out.
@@
First, I capture positional externalities using a feasibility condition: an interim status allocation is feasible (i.e., can be induced by an allocation of positional goods) if and only if it is weakly majorized by the buyers’ type distribution in the quantile space.
@@
Then, I focus on revenue maximization. Using results from Kleiner, Moldovanu and Strack (2021, henceforth KMS), I show that offering more levels of positional goods increases the seller’s revenue if the type distribution is regular in the sense of Myerson (1981). Therefore, the revenue-maximizing mechanism excludes types with negative marginal revenue and fully separates participants. This can be implemented by an all-pay auction with a reserve price, where the buyer who pays more obtains the higher position. In contrast, if Myerson’s regularity condition fails, the optimal mechanism will pool some types into the same level of positional goods.
I also find that a posted-price mechanism that offers a single good at a fixed price can guarantee at least half the maximum revenue. For uniform and exponential distributions, the approximation ratio exceeds 90%. The approximation result is relevant when the levels of positional goods is subject to regulatory or practical constraints. It also highlights the role of positional externalities: In the standard problem of selling indivisible ordinary goods, where there are no positional externalities, a posted-price mechanism is revenue-maximizing. By contrast, when selling positional goods, the seller cannot offer the highest status without reducing its value to all buyers who pay the posted price. Consequently, a posted-price mechanism is no longer optimal, and the seller’s revenue is strictly lower.
@@
Next, I study how the number of positional good levels and exclusion affect consumer welfare. I first focus on the case when consumer participation is fixed, for example, when there is a free regular service available. I show that the consumer surplus is higher when the seller offers more positional good levels if the buyer’s type distribution has an increasing failure rate (IFR), which implies a thin tail. This result highlights the tension between the seller’s revenue and the consumer surplus. On the other hand, when the type distribution has a decreasing failure rate (DFR), which implies a heavy tail, the consumer surplus is higher when the seller offers fewer positional good levels. Furthermore, under IFR, expanding consumer participation increases the consumer surplus, as the marginal gain of additional participants outweighs the inframarginal loss of the existing participants from the reduced status caused by pooling.
@@
I also study optimal mechanisms that maximize the consumer surplus. When negative transfers are allowed, full separation without exclusion maximizes the consumer surplus. Intuitively, this is because assortative matching is efficient, and possibly negative transfers can redistribute the efficiency gains among consumers and induce truth-telling (see also Gershkov and Schweinzer, 2010). By contrast, in applications where the price (i.e., effort) must be nonnegative, such as education and organizational hierarchies, I show that total pooling without exclusion maximizes the consumer surplus if the distribution satisfies IFR. Unlike selling ordinary goods, exclusion raises the status of participants who would be otherwise pooled with the excluded ones. Nevertheless, under IFR, the harm of exclusion to the excluded consumers exceeds its benefit to the participants. In contrast, if the distribution has a heavy tail (e.g., the Pareto distribution), the consumer surplus under the nonnegative price constraint is maximized by full separation without exclusion. Intuitively, separation benefits high types but hurts low types, so it benefits agents on average when there are relatively more high types.
@@
As an extension, I show that my main results extend to the setting where the intrinsic value of the positional good is determined by its intrinsic quality produced at a cost (à la Mussa and Rosen (1978)). Moreover, I also allow the seller to deliberately lower the status, or equivalently, prolong the waiting time by pausing service or procrastinating. Thus, the seller can create negative status or excessive waiting time as a substitute for exclusion. I show that the revenue-maximizing mechanism fully separates those with positive virtual valuations and assigns a negative status to those with negative virtual valuations. Additionally, my results extend to the setting where the status game is a “suffering” in that higher types suffer more from participation.
@@
The model has dual interpretations that lead to two strands of applications. On the one hand, it is a monopolist problem of selling positional goods, such as priority services (Gershkov and Winter, 2023) and luxury goods. Alternatively, it can be interpreted as an organizational design problem where the principal motivates agents’ effort (e.g., Moldovanu, Sela and Shi, 2007; Immorlica, Stoddard and Syrgkanis, 2015) or maximizes agents’ welfare through status within an organization. Under this interpretation, the optimal mechanism that maximizes effort possibly excludes some types and fully separates the participants if the ability distribution satisfies Myerson’s regularity. This can be seen in “rank-and-yank” organizations that rank every employee and terminate those who underperform.
@@
The implications also echo the education example by Hirsch (1976) and shed light on the consequences of education arms races. From the agent’s perspective, while the common wisdom advocates educational disarmament and less intense competition, I show that this benefits agents if their ability distribution has a thin tail: Under the IFR property, admission lotteries or random assignments, instead of meritocratic allocation, maximizes agents’ welfare. On the contrary, if a few agents have very high abilities (i.e., a heavy tail), meritocracy maximizes agents’ welfare on average. Intuitively, low-ability agents benefit from soft competition, while high-ability agents prefer meritocracy, so on average, meritocracy raises (reduces) agents’ welfare when there are relatively more high-ability (low-ability) agents. Additionally, if education is sufficiently socially productive, meritocracy can be optimal for social welfare.
@@
In Appendix A, I consider the case where exclusion is impossible, which is true in many applications where the lowest status is free. For example, a society may not allow exclusion, so investing no effort guarantees the lowest social status (Moldovanu, Sela and Shi, 2007). Moreover, priority services may be offered in addition to free regular services (Gershkov and Winter, 2023). In these cases, the lowest-level positional good may still have a positive value, as the lowest status cannot be offered to multiple agents at its original value either. I provide necessary and sufficient conditions when the revenue-maximizing mechanism is separating or pooling, and when consumer welfare is decreasing or increasing in the number of positional good levels. I also show some additional results, such as the 2-approximation of the maximum revenue by selling a single good above the lowest level. The results are qualitatively similar to the benchmark model with exclusion.
@@
The idea that the consumer’s utility depends on the comparison with others’ consumption dates back to Veblen (1899) (see also Duesenberry, 1949; Leibenstein, 1950; Bagwell and Bernheim, 1996). Since Hirsch (1976) introduced the concept of positional good, the status interdependence in preference has been formally modeled and studied by Frank (1985b, 2005, 2008), Robson (1992), and Hopkins and Kornienko (2004).  [^3] Another strand of literature uses the Duesenberry formulation that stems from Duesenberry (1949) and Pollak (1976) (see the survey by Truyts (2010)), which assumes the consumer’s utility depends on the consumption of others in addition to their absolute consumption. This literature focuses on consumer choice in the presence of positional externalities and the welfare effects of changes in the income distribution and government policies (e.g., luxury tax).
@@
Empirical evidence of positional goods and externalities include Frank (1985a, 2005), Luttmer (2005), Alpizar, Carlsson and Johansson-Stenman (2005), Carlsson, Johansson-Stenman and Martinsson (2007), Bursztyn et al. (2018). In particular, education is a typical example of positional good (Hirsch, 1976). Durst (2021) finds empirical evidence that education is indeed positional in Germany.
@@
Moldovanu, Sela and Shi (2007, henceforth MSS) study the optimal design of the number and size of status categories in an organization to maximize agents’ effort when they care about their relative position (i.e., status). They find that full separation is optimal if the ability distribution satisfies IFR, while coarse partitions may be optimal for some distributions—for example, the optimal partition involves only two categories if the distribution is sufficiently concave. Because status within an organization is also a positional good, my framework generalizes theirs by allowing for exclusion of agents and stochastic allocations, while further analyzing agents’ welfare.  [^4] The status definition in MSS—the number of strictly inferior agents minus the number of strictly superior agents—is equivalent (up to an affine transformation) to mine (see Footnote 10). In practice, organizations may exclude agents, and employee welfare is an important consideration. Methodologically, I use a mechanism design approach with a feasibility characterization using majorization, while they use methods in optimal contest design (with finite agents).  [^5] In the case without exclusion, I obtain stronger results that are necessary and sufficient conditions or require fewer distributional assumptions (see Appendix A).
@@
Relatedly, Rayo (2013) studies a model of monopolist provision of conspicuous goods and characterizes the optimal allocation. His definition of conspicuous goods differs from positional goods in that consumers care about the average type in their assigned category instead of relative positions. Nevertheless, his results are qualitatively similar to MSS’s, and he also focuses on revenue maximization with deterministic mechanisms and without exclusion. In Section 4.4, I also show that the feasibility condition can also be applied to conspicuous goods in his model, which explains the similarities between our results for revenue maximization.  [^6] Board (2009) assumes that the value of being in a group category takes a more general form and finds that the revenue-maximizing mechanism segregates agents too finely compared to the efficient level under a regularity condition. He also finds that revenue maximization excludes too many agents.
@@
Similar to MSS, Immorlica, Stoddard and Syrgkanis (2015) study the optimal design of badges to incentivize users’ contribution (i.e., effort) to the online community through the status value. Although they consider exclusion and allow the status value to be concave or convex in the relative position, they do not consider agents’ welfare or stochastic allocations either. Most importantly, their status value function is decreasing in the number of people weakly superior to them, without adjustment for those at the same level. This specification precludes pooling from being optimal because it reduces the status of all pooled agents. Consequently, total pooling is never optimal—even for agents’ welfare—in sharp contrast to my result, and exclusion is equivalent to pooling at the lowest status.
@@
In a different vein, Gershkov and Winter (2023, henceforth GW) study a model of priority service and analyze its implications for consumer welfare. Because priority service (PS) is a positional good arising from technological constraints, my model extends theirs to multiple (and stochastic) PS levels while allowing for exclusion using a mechanism design approach. By contrast, they focus on the welfare implications of having two PS levels instead of the optimal mechanism.  [^7] They extend their model to multiple PS levels, but without exclusion, in Section V. In the case without exclusion, I refine some of their results by providing necessary and sufficient conditions and show some additional results in Appendix A.
@@
Another closely related paper is by Loertscher and Muir (2022), who study revenue maximization when selling a fixed inventory of goods with heterogeneous qualities. Because positional externalities can arise from a limited stock, under the linear status specification and without the intrinsic value, my model can be viewed as a continuous version of theirs where each quality has a unit stock. In particular, pooling multiple types into the same status corresponds to randomization (or conflation) in their model. Nevertheless, my modeling approach leads to various different applications in positional goods, such as education and organizational status. Moreover, because positional externalities may also arise from psychological reasons, I also consider nonlinear (concave or convex) status specifications and different magnitudes of concern about the number of consumers at the same level.
@@
More broadly, this paper is also related to mechanism design with allocative externalities (e.g. Jehiel, Moldovanu and Stacchetti, 1996; Jehiel and Moldovanu, 2005; Csorba, 2008; Ostrizek and Sartori, 2023; Akbarpour, Dworczak and Kominers, 2024; Dworczak et al., 2024). In particular, this model focuses on a large population of agents, and the positional externalities arise from their relative consumption instead of types, group identities, or network effects.